What is the value of the y variable in the solution to the following system of equations?
3x + y = 14
–3x + 2y = 1

Answer by Maths68(1474)

3x + y = 14............(1)
–3x + 2y = 1 .........(2)
Add (1) and (2)
3x + y = 14
–3x + 2y = 1
-----------------
3y=15
3y/3=15/3
y=5